OpenStudy (anonymous):

how to solve log x * log 2

OpenStudy (anonymous):

Help is greatly appreciated, I'll have a test in about 20 more mins :) thanks.

OpenStudy (anonymous):

Technically, it's not solving since it's an expression, not an equation. ;-) Anyways, the property that you want to use here to simplify the expression is \(\log(ab) = \log(a)\log(b)\). Can you take things from here? :-)

OpenStudy (arthur_ser):

I don't think that it can be simplified any further at this form

OpenStudy (loser66):

is it not log a + log b = log (ab)? @ChristopherToni

OpenStudy (loser66):

@dalia_lam when you say "solve" , you need = sign. where is it? log(x) log (2) = ??? one more thing: log 2 is a constant. Therefore, to me, you just need = somenumber and solve it.

OpenStudy (anonymous):

@Loser66 That's what being tired does to you...sigh... XD You are indeed correct.

OpenStudy (loser66):

XD

OpenStudy (anonymous):



OpenStudy (anonymous):

it was the critical thinking question, to condense the expression to a single logarithm.